# Models That Prove Their Own Correctness

### v2

@@ Appendix B Preliminaries on Stochastic Gradient Ascent
Definition B.1 (Gradient estimator).
@@
1:Initialize $w_{0}\coloneqq\vec{0}\in\mathbb{R}^{d}$.
22:for $i=0,\dots,N-1$ do
3:Sample $v_{i}\sim D_{f}(w_{i-1})$.
3      4:Update $w_{i}\coloneqq w_{i-1}+\lambda\cdot v_{i}$.
5:Output $\bar{w}\coloneqq\frac{1}{N}\sum_{i\in[N]}w_{i}$.
@@
If at every iteration it holds that $||d_{i}||<B_{\mathrm{Lip}}$, then
@@
Appendix C Annotations
@@
In terms of efficiency, think of the annotator as an algorithm of the same computational resources as an honest prover in the system (see Definition 3.2, and the answer extractor as an extremely simple algorithm (e.g., trim a fixed amount of tokens from the annotation).
@@
Appendix D A simple proof system for the GCD
@@
Before we dive in, let us clarify what we mean by a proof system for the GCD. Paul has two integers $212$ and $159$; he claims that $GCD(212,159)=53$. An inefficient way for Veronica to check Paul’s answer is by executing the Euclidean algorithm on $(212,159)$ and confirm that the output is $53$. In an efficient proof system, Veronica asks Paul for a short string $\pi^{\ast}$ (describing two integers) with which she can easily compute the answer—without having to repeat Paul’s work all over. On the other hand, if Paul were to claim that “$GCD(212,159)=51$” (it does not), then for any alleged proof $\pi$, Veronica would detect an error and reject Paul’s claim.
@@
Claim D.1 (Bézout’s identity [Bezout, 1779]).
@@
Any coefficients $z_{0},z_{1}$ satisfying the assumption of D.1 are known as Bézout coefficients for $(x_{0},x_{1})$. D.1 immediately gives our simple proof system: For input $x=(x_{0},x_{1})$ and alleged GCD $y$, the honest prover sends (alleged) Bézout coefficients $(z_{0},z_{1})$. The Verifier accepts if and only if $y=z_{0}\cdot x_{0}+z_{1}\cdot x_{1}$ and $y$ divides both $x_{0}$ and $x_{1}$.
In this proof system the Verifier does not need to make any query; to fit within Definition 3.2, we can have the verifier issue a dummy query. Furthermore, by D.1 it is complete and has soundness error $s=0$. Lastly, we note that the Verifier only needs to perform two multiplications, an addition, and two modulus operations; in that sense, verification is more efficient than computing the GCD in the Euclidean algorithm as required by Remark 3.3.
@@
To describe how a proof $z=(z_{0},z_{1})$ is annotated, let us first note how it can be computed. The Bézout coefficients can be found by an extension of the Euclidean algorithm. It is described in Algorithm 4.999Our description is the same as https://en.wikipedia.org/wiki/Extended_Euclidean_algorithm.
@@
1:Initialize $r_{0}=x_{0}$, $r_{1}=x_{1}$, $s_{0}=1$, $s_{1}=0$, and $q=0$.
12:while $r_{1}\neq 0$ do
3:Update $q\coloneqq\lfloor r_{0}/r_{1}\rfloor$.
4:Update $(r_{0},r_{1})\coloneqq(r_{1},r_{0}-q\times r_{1})$.
2      5:Update $(s_{0},s_{1})\coloneqq(s_{1},s_{0}-q\times s_{1})$.
6:Output GCD $y=r_{0}$ and Bézout coefficients $z_{0}\coloneqq s_{0}$ and $z_{1}\coloneqq(r_{0}-s_{0}\cdot x_{0})/x_{1}$.
@@
We formalize this in the terminology of Appendix C by defining a Transcript Annotator and Answer Extractor. Note that, since our proof system consists only of one “answer” $z$ send from the prover to the verifier, the entire transcript $\pi$ is simply $z=(z_{0},z_{1})$. Since the verification is deterministic, this means that the proof system is of an $\mathsf{NP}$ type (however, note that the search problem of finding the “$\mathsf{NP}$-witness” $z=(z_{0},z_{1})$ is in fact in $\mathsf{P}$).
@@
Appendix E Experiment details
We provide details of how we implemented the experiments in Section 5 and additional figures for each experiment.
@@ Appendix E Experiment details
We use Karpathy’s nanoGPT. [^11] https://github.com/karpathy/nanoGPT. We use a 6.7M parameter architecture of $8$ layers, $8$ attention heads, and $256$ embedding dimensions. We optimized hyperparameters via a random hyperparameter search, arriving at learning rate $0.0007$, AdamW $\beta_{1}=0.733$ and $\beta 2=0.95$, $10\%$ learning rate decay factor, no dropout, gradient clipping at $2.0$, no warmup iterations, and $10\%$ weight decay.
@@
We sample integers from the $\log_{10}$-uniform distribution over $\{1,\dots,10^{4}\}$. Models in Tables 2 and 2 are trained for 100K iterations on a dataset of of $\approx$10M samples. For Figure 3 (base ablation) we train for 20K iterations on a dataset of $\approx$1M samples; this is because this setting required 68 many runs in total, whereas the annotation-cutoff ablation required 18 longer runs.
@@
All experiments were run on a machine with an NVIDIA A10G GPU, 64GB of RAM, and 32 CPU cores. Longer runs (annotation-cutoff ablation) took about 75 minutes each. Shorter runs (base ablation) took about 15 minutes. The total running time of our experiments was approximately 40 hours, excluding time dedicated to a random hyperparameter search. The overall disk space needed for our models and data (to be made available upon publication) is 4GB.
@@
Annotations are encoded as above, with each component in an intermediate step $\pi_{t}$ delimited by a unique token. Since different integer pairs may require a different number of intermediate steps to compute the Bézout coefficients, we chose to pad all annotaitons to the same length $T$ by the last step $\pi_{T}$ in the sequence (which consists of the final Bézout coefficients). This ensures that the final component output by the model in each sequence should be the Bézout coefficient, and allows us to batch model testing (generation and evaluation) resulting in a 1000x speed-up over sequential testing.

### v3

@@ Appendix C Preliminaries on Stochastic Gradient Ascent
Definition C.1 (Gradient estimator).
@@
2Initialize $w_{0}\coloneqq\vec{0}\in\mathbb{R}^{d}$.
3 for $i=1,\dots,N-1$ do
4       Sample $v_{i}\sim D_{f}(w_{i-1})$.
5       Update $w_{i}\coloneqq w_{i-1}+\lambda\cdot v_{i}$.
6
Output $\bar{w}\coloneqq\frac{1}{N}\sum_{i\in[N]}w_{i}$.
@@
If at every iteration it holds that $||v_{i}||<B_{\mathrm{Lip}}$, then
@@
C.1 Learning with Stochastic Gradient Ascent/Descent
C.2 captures the general case of using SGA for maximization of concave problems. It is more common for the literature to discuss the equivalent setting of Stochastic Gradient Descent (SGD) for minimization of convex problems. Specifically, a common application of SGD is for the task of Risk Minimization: given a loss function and access to an unknown distribution of inputs, the goal is to minimize the expected loss with respect to the distribution. Assuming that the loss function is differentiable, the gradient of the loss serves as a gradient estimator (see Definition C.1) for the risk function. We refer the reader to Shalev-Shwartz and Ben-David [2014, Section 14.5.1] for a complete overview of SGD for risk minimization.
For the sake of completeness, we formulate Transcript Learning (TL, Algorithm 1) in the framework of Risk Minimization for Supervised Learning. Although multiple loss functions may achieve our ultimate goal—learning Self-Proving models—in what follows we define the loss that corresponds to TL. Fix a verifier $V$ and let $\mathcal{T}^{\ast}_{V}$ denote a distribution over accepting transcripts. We define
$\textsf{loss}\left(\theta,(x,\pi^{\ast})\right)\coloneqq\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}(x)}\left[\pi\neq\pi^{\ast}\right],$ (10)
where $\pi^{\ast}$ and $\pi$ share the same verifier messages (as in Lemma B.4) so the inequality is only over the prover’s messages, namely $\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}(x)}\left[\pi\neq\pi^{\ast}\right]=\Pr_{\pi\sim\mathcal{T}^{\theta}_{V}(x)}[(y,a_{1},\dots,a_{R})\neq(y^{\ast},a_{1}^{\ast},\dots,a_{R}^{\ast})]$. [^11] This loss is not to be confused with those discussed in Appendix A. Here, we are simply explaining how TL can be viewed as a supervised risk minimizer for the loss function defined in Equation 10.
The risk function is the expected value of the loss over the joint distribution of inputs and accepting transcripts $\mu\times\mathcal{T}^{\ast}_{V}(\mu)$:
$\textsf{Risk}\left(\theta\right)\coloneqq\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ \pi^{\ast}\sim\mathcal{T}^{\ast}_{V}\end{subarray}}\left[\textsf{loss}\left(\theta,(x,\pi^{*})\right)\right],$
which means that the agreement function defined in Theorem B.5
$A(\theta)=\Pr_{\begin{subarray}{c}x\sim\mu\\ \pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)\\ \pi\sim\mathcal{T}^{\theta}_{V}(x)\end{subarray}}\left[\pi=\pi^{*}\right]$
satisfies $A(\theta)=1-\textsf{Risk}(\theta)$.
Thus, maximizing the agreement is equivalent to minimizing the risk. The hypothesis class over which the optimization is performed is the ball of radius $B_{\mathrm{Norm}}$, i.e., $\left\{\theta\in\mathbb{R}^{d}:||\theta||<B_{\mathrm{Norm}}\right\}$. The assumption that $A$ is concave in $\theta$ implies that the loss function is convex in $\theta$, which is the required assumption for using SGD for risk minimization.
Indeed, TL uses the natural gradient estimator for this setting, the gradient of the “complement” of the loss: $\Pr_{\pi}\left[\pi=\pi^{\ast}\right]$, since TL maximizes the agreement instead of minimizing the risk. The proof of Lemma B.4, i.e., $\nabla_{\theta}A(\theta)=\operatorname*{\mathbb{E}}_{x,\pi^{\ast}}\left[\nabla_{\theta}\left(\Pr_{\pi}\left[\pi=\pi^{\ast}\right]\right)\right]$, follows from the above discussion.
Appendix D Annotations
@@
In terms of efficiency, think of the annotator as an algorithm of the same computational resources as an honest prover in the system (see Definition 3.2), and the answer extractor as an extremely simple algorithm (e.g., trim a fixed amount of tokens from the annotation).
@@
Appendix E A simple proof system for the GCD
@@
Before we dive in, let us clarify what we mean by a proof system for the GCD. Prover Paul has two integers $212$ and $159$; he claims that $GCD(212,159)=53$. An inefficient way for Verifier Veronica to check Paul’s answer is by executing the Euclidean algorithm on $(212,159)$ and confirm that the output is $53$. In an efficient proof system, Veronica asks Paul for a short string $\pi^{\ast}$ (describing two integers) with which she can easily compute the answer—without having to repeat Paul’s work all over. On the other hand, if Paul were to claim that “$GCD(212,159)=51$” (it does not), then for any alleged proof $\pi$, Veronica would detect an error and reject Paul’s claim.
@@
Claim E.1 (Bézout’s identity [Bezout, 1779]).
@@
Any coefficients $z_{0},z_{1}$ satisfying the assumption of E.1 are known as Bézout coefficients for $(x_{0},x_{1})$. E.1 immediately gives our simple proof system: For input $x=(x_{0},x_{1})$ and alleged GCD $y$, the honest prover sends (alleged) Bézout coefficients $(z_{0},z_{1})$. The Verifier accepts if and only if $y=z_{0}\cdot x_{0}+z_{1}\cdot x_{1}$ and $y$ divides both $x_{0}$ and $x_{1}$.
In this proof system the Verifier does not need to make any query; to fit within Definition 3.2, we can have the verifier issue a dummy query. Furthermore, by E.1 it is complete and has soundness error $s=0$. Lastly, we note that the Verifier only needs to perform two multiplications, an addition, and two modulus operations; in that sense, verification is more efficient than computing the GCD in the Euclidean algorithm as required by Remark 3.3.
@@
To describe how a proof $z=(z_{0},z_{1})$ is annotated, let us first note how it can be computed. The Bézout coefficients can be found by an extension of the Euclidean algorithm. It is described in Algorithm 4.121212Our description follows https://en.wikipedia.org/wiki/Extended_Euclidean_algorithm.
@@
1 Initialize $r_{0}=x_{0}$, $r_{1}=x_{1}$, $s_{0}=1$, $s_{1}=0$, and $q=0$.
2 while $r_{1}\neq 0$ do
3       Update $q\coloneqq\lfloor r_{0}/r_{1}\rfloor$.
4       Update $(r_{0},r_{1})\coloneqq(r_{1},r_{0}-q\times r_{1})$.
5       Update $(s_{0},s_{1})\coloneqq(s_{1},s_{0}-q\times s_{1})$.
6
Output GCD $y=r_{0}$ and Bézout coefficients $z_{0}\coloneqq s_{0}$ and $z_{1}\coloneqq(r_{0}-s_{0}\cdot x_{0})/x_{1}$.
@@
We formalize this in the terminology of Appendix D by defining a Transcript Annotator and Answer Extractor. Note that, since our proof system consists only of one “answer” $z$ send from the prover to the verifier, the entire transcript $\pi$ is simply $z=(z_{0},z_{1})$. Since the verification is deterministic, this means that the proof system is of an $\mathsf{NP}$ type (however, note that the search problem of finding the “$\mathsf{NP}$-witness” $z=(z_{0},z_{1})$ is in fact in $\mathsf{P}$).
@@
Appendix F Experiment details
We provide details of how we implemented the experiments in Section 5 and additional figures for each experiment. Code, data and models are available at https://github.com/orrp/self-proving-models.
@@ Appendix F Experiment details
We use Karpathy’s nanoGPT [^14] https://github.com/karpathy/nanoGPT. implementation of GPT. Note that we train the model “from scratch” only on sequences related to the GCD problem, rather than starting from a pretrained checkpoint. We use a 6.3M parameter architecture of $8$ layers, $8$ attention heads, and $256$ embedding dimensions. We optimized hyperparameters via a random hyperparameter search, arriving at learning rate $0.0007$, AdamW $\beta_{1}=0.733$ and $\beta 2=0.95$, $10\%$ learning rate decay factor, no dropout, gradient clipping at $2.0$, no warmup iterations, and $10\%$ weight decay.
@@
We sample integers from the $\log_{10}$-uniform distribution over $\{1,\dots,10^{4}\}$. Models in Tables 2 and 2 are trained for 100K iterations on a dataset of $\approx$10M samples. For Figure 3 (base ablation) we train for 20K iterations on a dataset of $\approx$1M samples; this is because this setting required 68 many runs in total, whereas the annotation-cutoff ablation required 18 longer runs.
@@
All experiments were run on a machine with an NVIDIA A10G GPU, 64GB of RAM, and 32 CPU cores. The longest experiment was the single RLVF run, which took one month and four days. The annotation-cutoff ablation runs took about 75 minutes each. Base of representation ablation runs were shorter at about 15 minutes each. The total running time of the Transcript Learning experiments was approximately 40 hours (excluding time dedicated to a random hyperparameter search), and the RLVF experiment took another month and four days. The overall disk space needed for our models and data is 4GB.
@@
Annotations are encoded as above, with each component in an intermediate step $\pi_{t}$ delimited by a unique token. Since different integer pairs may require a different number of intermediate steps to compute the Bézout coefficients, we chose to pad all annotations to the same length $T$ by the last step $\pi_{T}$ in the sequence (which consists of the final Bézout coefficients). This ensures that the final component output by the model in each sequence should be the Bézout coefficient, and allows us to batch model testing (generation and evaluation) resulting in a 1000x speed-up over sequential testing.
@@
Figure 6: RLVF Verifiability as a function of the number of samples $N$. Starting from a base model with Verifiability 48% (obtained via Transcript Learning), in each iteration a batch of 2048 inputs are sampled; the model generates a proof for each; the Verifier is used to check which proofs are accepted; then, the model parameters are updated accordingly (see Algorithm 2). Verifiability was evaluated on a held-out dataset of 1k inputs.